Ty's scores on four test are 83,91,78, and 81. How many points above the average must Ty score on the next test in order to raise his average 2 points?
Answer by Cromlix(4381)

Hi there,
Totalling 83,91, 78 and 81 = 333
Average = 333/4 = 83.25
To score two points = 85.25
5 tests
5 x 85.25 = 426.25
So, Ty would have to score
426.25 - 333 = 93.25
Hope this helps:-)